A sample of metal weighs 210 g in air , 180 g in water and 120 g i an unknown liquid .then what will be density of that metal?

Grade:9

6 Answers

surudhi
8 years ago
The relative density of a metal = Weight of the metal in air / Loss in weight of metal in water = 210/(210-180) = 210/30 = 7.
anonymous
7 years ago
relative density = (weight of body in air -weight of body in liquid)/(weight of the body in air – weight of body in water)
weight of the body in air  =210g
weight of body in water = 180g
weight of body in liquid = 120g
 
(210-120) / (210-180) = 3
 
relative density = 3
Jdlsjdfhj
7 years ago
Metal=210/(210-180)= 7
In the Same way
Relative density of liquid=(210-120)/(210-180)=90/30=3
Ans) 7,3
Sex
Nitish
6 years ago
Hello my name is Nitish I will give you the best answer The answer is you can come again and again in my house I will try my best to give the answer and then we can do enjoy by doing lunch and then we can go to the big hotel and can eat the dinner with me and please come girls only then I will try to make sure to be the best student of the class
Shifts
5 years ago
Do Ken of to it to it so he do it to it we it we it we have this Ken push me to it to felon man if this funk Ken fuc Dennis rush named thenthanks K
Yash Chourasiya
3 years ago
Dear Student

Relative density of metal = wieght in air / Change in weight of water
= (210) / (210 – 180)
Denstiy of metal = 7 g/cm2

I hope this answer will help you.
Thanks & Regards
Yash Chourasiya